OpenStudy (anonymous):

Using laws of exponents, explain how to simplify (x-2)3

OpenStudy (angela210793):

=(x-3)*(x-3)^2=(x-3)*(x^2-6x+9)

OpenStudy (anonymous):

(x-2)^3 = (x-2)(x-2)(x-2)=(x^2-4x+4)(x-2) =(x^3-4x^2+4x-2x^2+8x-8) = (x^3 -6x^2+12x-8)